# Evaluating Memory in LLM Agents via Incremental Multi-Turn Interactions

### v1

@@ A.1 Accurate Retrieval (AR)
This is a dialogue-based QA dataset. For LME(S*), we use multiple historical conversation data segments, arrange them in chronological order, and finally concatenate them to create five long conversation histories, each with a length of approximately 355K tokens. Since some of the questions have open-ended answers, we adopt the approach used in previous work and employ the GPT-4o model to assess whether the agent’s responses meet the requirements. If a response is deemed satisfactory, it is marked as True. Finally, we calculate the proportion of satisfactory responses as the evaluation metric.
@@
Using five books from Bench (each >390K tokens, counted using the gpt-4o-mini tokenizer), we identify the ten most frequently mentioned characters via SpaCy NER. We extract 101 events experienced by key characters using gpt-4o. For each event, we construct a 6-way multiple-choice question by pairing the true event with five distractors generated via gpt-4o. The agent receives up-to five previous events and must identify the correct continuation. We report the mean accuracy over 100 such questions per book, and ultimately present the average accuracy across all five books.
| Category | Dataset | Metrics | Avg. Length | Description |
| Accurate Retrieval | RULER-QA1 | SubEM | 197K | Gold passage retrieval QA. |
| RULER-QA2 | 421K | |||
| RULER-NIAH-MQ | Recall | 448K | Retrieve multiple “needles” from the “haystack”. | |
| Bench-QA | ROUGE F1 | 183K | Novel QA with entity replacement. | |
| LongMemEval (S) | Model Based Acc. | 110K | Dialogues based QA. | |
| LongMemEval (S*) | 355K | |||
| EventQA (ours) | Accuracy | 534K | Novel multiple-choice QA on characters events. | |
| Test-time Learning | BANKING77 | Accuracy | 103K | Banking intent classification, 77 labels |
| CLINC150 | Intent classification, 151 labels | |||
| NLU | Task intent classification, 68 labels | |||
| TREC Coarse | Question type classification, 6 labels | |||
| TREC Fine | Question type classification, 50 labels | |||
| Long Range Understanding | ||||
| Bench-Sum | Model Based F1 | 172K | Novel summarization with entity replacement. | |
| Conflict Resolving | FactConsolidation-SH (ours) | SubEM | 262K | Conflict solving in single hop reasoning. |
| FactConsolidation-MH (ours) | Conflict solving in multiple hop reasoning. |
@@ A.3 Long-Range Understanding (LRU)
We evaluate LRU via the Summarization task En.Sum from -Bench [53]. We follow the settings from [50] and use the GPT-4o model in evaluating the summarized text. In this process, we assess the fluency of the input text (scored as 0 or 1) and use the dot product of this score with the F1 score as the final evaluation metric.
A.4 Conflict Resolution (CR)
@@
We use counterfactual edit pairs from the MQUAKE [54] dataset. Each sentence containing information was assigned a number. For each edit pair, the sentence representing outdated information (the distractor) is given a smaller number, while the sentence representing more recent information (the one containing the answer) is given a larger number. We then concatenate these sentences into a long context in order according to their assigned numbers. We evaluate the CR via two datasets: Single-Hop Editing and Multi-Hop Editing. In these tasks, the agent’s responses are mostly informational entities. Therefore, we also use SubEM (Substring Exact Match) as the evaluation metric.
Appendix B Prompts
We introduce some example prompts used in this section.
B.1 Instructions for Memory Construction
@@
B.2 Instructions for Long-Context Agents
In Figure 5, we provide the examples of instructions used on different of datasets. For some existing datasets, we adjust the prompt settings from previous work such as [15, 48]. For example, for the dataset Bench-QA and Bench-Sum, we also insert two answer examples as <demo> in the prompt to help the agent better understand the questions and standardize its outputs.
B.3 Instructions for RAG Agents
We provide examples of prompts used for the RAG based Agents in Figure 6. For this type of agent, after storing the input long context in memory, we use <question> as the memory retrieval query for most tasks. But for task RULER-NIAH-MQ, we use the entire question "What are all the special magic numbers for <question> mentioned in the memorized text?" as the query. And for Bench-Sum, we use the entire query without the <demo> for the memory retrieval.
| Agent Type | RULER-QA1 | RULER-QA2 | NIAH-MQ | Bench-QA | LME(S) | LME(S*) | EventQA |
| Long-Context Agents | |||||||
| GPT-4o | 72.0 | 51.0 | 25.0 | 55.4 | 61.4 | 32.0 | 77.2 |
| GPT-4o-mini | 64.0 | 43.0 | 22.8 | 44.9 | 55.6 | 30.7 | 59.0 |
| GPT-4.1-mini | 83.0 | 66.0 | 94.8 | 45.8 | 61.4 | 55.7 | 82.6 |
| Gemini-2.0-Flash | 87.0 | 59.0 | 83.8 | 53.2 | 52.6 | 47.0 | 67.2 |
| Claude-3.7-Sonnet | 77.0 | 53.0 | 38.0 | 50.6 | 59.0 | 34.0 | 74.6 |
| GPT-4o-mini | 64.0 | 43.0 | 22.8 | 44.9 | 55.6 | 30.7 | 59.0 |
| Simple RAG Agents | |||||||
| BM25 | 68.0 | 54.0 | 100.0 | 45.6 | 55.2 | 45.3 | 74.6 |
| Embedding RAG Agents | |||||||
| Contriever | 22.0 | 31.0 | 2.5 | 38.1 | 32.8 | 15.7 | 66.8 |
| Text-Embed-3-Small | 60.0 | 44.0 | 7.2 | 44.4 | 49.0 | 48.3 | 63.0 |
| Text-Embed-3-Large | 54.0 | 44.0 | 19.5 | 50.1 | 44.6 | 52.3 | 70.0 |
| NV-Embed-v2 | 90.0 | 67.0 | 73.5 | 51.4 | 45.4 | 55.0 | 72.8 |
| Structure-Augmented RAG Agents | |||||||
| RAPTOR | 29.0 | 38.0 | 15.8 | 31.3 | 38.8 | 34.3 | 45.8 |
| GraphRAG | 47.0 | 47.0 | 38.3 | 35.8 | 39.2 | 35.0 | 34.4 |
| HippoRAG-v2 | 76.0 | 66.0 | 67.5 | 45.7 | 44.2 | 50.7 | 67.6 |
| Mem0 | 24.0 | 32.0 | 4.8 | 22.4 | 45.0 | 36.0 | 37.5 |
| Cognee | 31.0 | 26.0 | 4.0 | 19.7 | 31.3 | 29.3 | 26.8 |
| Agentic Memory Agents | |||||||
| Self-RAG | 35.0 | 42.0 | 8.0 | 28.5 | 23.4 | 25.7 | 31.8 |
| MemGPT | 41.0 | 38.0 | 8.8 | 20.8 | 41.4 | 32.0 | 26.2 |
| Agent Type | BANKING | CLINC | NLU | TREC C | TREC F | Recom | Bench-Summ | FactCon-SH | FactCon-MH |
| Long-Context Agents | |||||||||
| GPT-4o | 96.0 | 96.0 | 90.0 | 87.0 | 69.0 | 12.3 | 32.2 | 60.0 | 5.0 |
| GPT-4o-mini | 93.0 | 93.0 | 87.0 | 73.0 | 66.0 | 15.1 | 28.9 | 45.0 | 5.0 |
| GPT-4.1-mini | 93.0 | 82.0 | 85.0 | 68.0 | 50.0 | 16.7 | 41.9 | 36.0 | 5.0 |
| Gemini-2.0-Flash | 91.0 | 90.0 | 84.0 | 88.0 | 67.0 | 8.7 | 23.9 | 30.0 | 3.0 |
| Claude-3.7-Sonnet | 97.0 | 98.0 | 86.0 | 87.0 | 79.0 | 18.3 | 52.5 | 43.0 | 2.0 |
| GPT-4o-mini | 93.0 | 93.0 | 87.0 | 73.0 | 66.0 | 15.1 | 28.9 | 45.0 | 5.0 |
| Simple RAG Agents | |||||||||
| BM25 | 89.0 | 89.0 | 84.0 | 62.0 | 53.0 | 13.6 | 20.9 | 56.0 | 3.0 |
| Embedding RAG Agents | |||||||||
| Contriever | 89.0 | 88.0 | 80.0 | 55.0 | 41.0 | 15.2 | 21.2 | 18.0 | 7.0 |
| Text-Embed-3-Small | 88.0 | 89.0 | 83.0 | 54.0 | 36.0 | 15.3 | 25.7 | 28.0 | 3.0 |
| Text-Embed-3-Large | 90.0 | 91.0 | 80.0 | 55.0 | 46.0 | 16.2 | 21.6 | 28.0 | 4.0 |
| NV-Embed-v2 | 88.0 | 89.0 | 82.0 | 40.0 | 48.0 | 13.5 | 20.7 | 55.0 | 6.0 |
| Structure-Augmented RAG Agents | |||||||||
| RAPTOR | 78.0 | 75.0 | 73.0 | 48.0 | 23.0 | 12.3 | 13.4 | 14.0 | 1.0 |
| GraphRAG | 64.0 | 54.0 | 49.0 | 24.0 | 6.0 | 9.8 | 0.4 | 14.0 | 2.0 |
| HippoRAG-v2 | 81.0 | 86.0 | 73.0 | 38.0 | 29.0 | 10.2 | 14.6 | 54.0 | 5.0 |
| Mem0 | 5.0 | 4.0 | 1.0 | 6.0 | 1.0 | 10.0 | 0.8 | 18.0 | 2.0 |
| Cognee | 34.0 | 42.0 | 42.0 | 41.0 | 18.0 | 10.1 | 2.3 | 28.0 | 3.0 |
| Agentic Memory Agents | |||||||||
| Self-RAG | 19.0 | 13.0 | 6.0 | 15.0 | 5.0 | 12.8 | 0.9 | 19.0 | 3.0 |
| MemGPT | 89.0 | 83.0 | 79.0 | 56.0 | 31.0 | 14.0 | 2.5 | 28.0 | 3.0 |
Appendix C Detailed Experimental Results
@@
C.1 Detailed Results on AR
In Table 7, we present the detailed results for each agent on every dataset. For AR tasks, using Simple RAG Agents equipped with retrievers like BM25 can significantly improve performance compared to the backbone model. This is because the GPT-4o-mini is limited by its 128K context length, which restricts the amount of information it can process at once. Meanwhile, the overall performance of Embedding RAG Agents surpasses that of both Structure-Augmented RAG Agents and Agentic Memory Agents. This advantage is primarily attributed to the use of dense retrieval in Embedding RAG Agents. It enables the extraction of longer contextual information from memory. As a result, Embedding RAG Agents are able to provide richer and more comprehensive context for tasks.
C.2 Detailed Results on TTL, LRU and CR
We give detailed results on each dataset in Table 7. For all three types of tasks, RAG-based agents generally underperform compared to their respective GPT-4o-mini backbones. This observation highlights certain limitations inherent to the RAG approach. For instance, in TTL tasks, RAG-based methods often struggle to accurately retrieve context from memory that is closely associated with the input. In LRU tasks, these methods face challenges in achieving a comprehensive understanding of long contexts. Furthermore, for CR tasks—especially the multi-hop variants—effective handling requires strong reasoning and information extraction capabilities, which remain beyond the reach of most current agents.
C.3 Detailed Results on Ablation Study
In this section, we introduce the detailed results on the ablation study on different chunk sizes, retrieve number, context length and computation latency.
In Table 8 and 11, we compare the RAG-based Agents on different chunk sizes and datasets. We selected chunk sizes from the two sets {512, 4096} and {512, 1024, 2048, 4096}. For some datasets composed of synthetic text, such as RULER-QA, using a smaller chunk size generally helps RAG Agents and Agentic Memory Agents achieve better test performance. However, for datasets composed of continuous text, such as Bench-QA, since the retrieval number remains unchanged, reducing the chunk size does not lead to performance improvement.
In Table 11, we evaluate the selected RAG-based Agents on five datasets. We choose different TopK ranging from {2, 5, 10}. We find that for the AR series of tasks, increasing the retrieve number (TopK) leads to a significant improvement in performance. However, for the TTL series of tasks, the performance gains from increasing TopK are much less pronounced.

### v2

@@ A.1.2 Details on AR datasets
This is a dialogue-based QA dataset. For LME(S*), we use multiple historical conversation data segments, arrange them in chronological order, and finally concatenate them to create five long conversation histories, each with a length of approximately 355K tokens. Since some of the questions have open-ended answers, we adopt the approach used in previous work and employ the GPT-4o model to assess whether the agent’s responses meet the requirements. If a response is deemed satisfactory, it is marked as True. Finally, we calculate the proportion of satisfactory responses as the evaluation metric. Wu et al. (2025) reported in Table 6 that a prompt-engineered GPT-4o judge achieves 98.0% accuracy and demonstrates very high stability.
@@
Using five books from -Bench (each contains more than 390K tokens, counted using the gpt-4o-mini tokenizer), we identify the ten most frequently mentioned characters via SpaCy NER. We extract 101 events experienced by key characters using gpt-4o. For each event, we construct a 6-way multiple-choice question by pairing the true event with five distractors generated via gpt-4o. The agent receives up-to five previous events and must identify the correct continuation. We report the mean accuracy over 100 such questions per book, and ultimately present the average accuracy across all five books.
@@ A.2 Test-time Learning (TTL)
A.2.1 Definition of TTL
A.2.2 Details on TTL datasets
We evaluate TTL via two task categories:
(1) Multi-Class Classification (MCC)
@@ A.2.2 Details on TTL datasets
(2) Recommendation (Recom.)
We concatenate multiple short dialogues about movie recommendations from the original dataset, remove duplicate dialogues, and create a long context containing over a thousand recommendation instances. In this task, the agent is required to recommend 20 movies based on the content of the dialogue. We evaluate the recommendations by calculating Recall@5, which measures the overlap between the top 5 recommended movies and the ground truth.
@@ A.3.2 Details on LRU datasets
We evaluate LRU via the Summarization task En.Sum from -Bench (Zhang et al., 2024). We follow the settings from (Yen et al., 2024) and use the GPT-4o model in evaluating the summarized text. In this process, we assess the fluency of the input text (scored as 0 or 1) and use the dot product of this score with the F1 score as the final evaluation metric.
A.4 Selective Forgetting (SF)
A.4.1 Definition of SF
In long-term interactions, agents often face evolving or conflicting information—whether about the external world (e.g., changes in political leadership) or user-specific facts (e.g., a new occupation). This challenge is closely related to model editing (Meng et al., 2023; Fang et al., 2024) and knowledge unlearning (Wang et al., 2024e), which focus on modifying or removing factual knowledge from language models. We define Selective Forgetting (SF) as the agent’s ability to detect and resolve contradictions between out of date knowledge and newly acquired information, ensuring the agent remains aligned with current realities and user states. SF is distinct from Abstractive Retrieval (AR) in two key ways. (1) Certain questions requiring SF cannot be answered solely through AR. As illustrated in Figure 1, an agent that retrieves all facts related to pears may fail to identify the updated information in the second message. (2) In AR, earlier messages remain relevant and should be retained, even when multiple pieces of evidence are required. In contrast, SF involves identifying outdated or incorrect information and discarding it. That is, AR requires preservation of all related content, whereas SF requires overwriting prior facts to reflect the most up-to-date truth.
A.4.2 Details on SF datasets
@@
We use counterfactual edit pairs from the MQUAKE (Zhong et al., 2023) dataset. Each sentence containing information was assigned a number. For each edit pair, the sentence representing outdated information (the distractor) is given a smaller number, while the sentence representing more recent information (the one containing the answer) is given a larger number. We then concatenate these sentences into a long context in order according to their assigned numbers. We evaluate the SF via two datasets: Single-Hop FactConsolidation and Multi-Hop FactConsolidation. In these tasks, the agent’s responses are mostly informational entities. Therefore, we also use SubEM (Substring Exact Match) as the evaluation metric to calculate the accuracy of QA.
A.5 Justification for competencies based on cognitive science
Accurate retrieval is central to human memory research, as evidenced by classical forgetting curves and recall tests that foreground fidelity of recall (Ebbinghaus, 2013). However, a sole focus on accuracy obscures another fundamental axis: the timescale of learning and consolidation. Ebbinghaus observed that an initial, fleeting grasp rarely endures without reinforcement (Ebbinghaus, 2013), and James (1890) distinguished primary (immediate) from secondary (enduring) memory. These classic distinctions ground our notions of test-time learning (incorporation of new information via memory) and long-range understanding (durable, integrated knowledge). Consistent with this, the Complementary Learning Systems (CLS) framework delineates a hippocampal system for rapid episodic learning and a neocortical system for gradual, structured knowledge accumulation, underscoring the need to assess both quick memorization and long-horizon retention (McClelland et al., 1995).
Beyond the acquisition–consolidation axis, another equally fundamental challenge is selective forgetting. Overlapping or contradictory traces can impede retrieval, and interference has long been recognized as a primary driver of forgetting in cognitive psychology (Anderson & Neely, 1996). Neurocognitive evidence further shows that the brain engages targeted control mechanisms to resolve such interference at retrieval time (Wimber et al., 2015). We therefore include selective forgetting—the ability to handle interference and contradictions—as a core dimension.
In sum, our four categories—accurate retrieval, test-time learning, long-range understanding, and selective forgetting—align with key dimensions of memory identified in cognitive science and AI memory systems, covering the essential capabilities that any robust memory mechanism must support in practice.
| Capability | Tasks | # of Sequences : QAs | Avg Len |
| Accurate Retrieval | SH-Doc QA | 1 : 100 | 197K |
| MH-Doc QA | 1 : 100 | 421K | |
| LongMemEval (S*) | 5 : 300 | 355K | |
| EventQA | 5 : 500 | 534K | |
| Test-Time Learning | BANKING-77 | 1 : 100 | |
| CLINC-150 | 1 : 100 | ||
| NLU | 1 : 100 | 103K | |
| TREC (Coarse) | 1 : 100 | ||
| TREC (Fine) | 1 : 100 | ||
| Movie-Rec Redial | 1 : 200 | 1.44M | |
| Long-Range Understanding | Bench-Sum | 100 : 100 | 172K |
| Detective QA | 10 : 71 | 124K | |
| Selective Forgetting | FactConsolidation-SH | 1 : 100 | 262K |
| FactConsolidation-MH | 1 : 100 |
Appendix B Detailed Memory Agents Description
We give detailed description of the memory agents used in experiments in this section.
B.1 Long-Context Agents
We evaluate five modern long-context LLMs: GPT-4o (OpenAI, 2025b) serves as the high-performance, low-latency model with better cost efficiency than prior generations. While GPT-4o-mini is a lightweight, budget-friendly alternative that enables large-scale evaluations by delivering faster responses and lower per-token costs. Notably, the GPT-4.1 (OpenAI, 2025a) family strengthens instruction following and maintains strong performance at very large context windows (reported up to 1M tokens). Considering the higher token cost, we choose the GPT-4.1-mini in evaluation. Gemini-2.0-Flash (DeepMind, 2025) targets high throughput and the use of built-in tools, offering a 1M token context window for efficient long-context processing. Claude-3.7-Sonnet (Anthropic, 2025) is a hybrid-reasoning model with optional visible “extended thinking,” strong math/coding skills, and developer-controlled thinking budgets.
B.2 RAG Agents
We consider three RAG variants: Simple RAG Agents, Embedding-based RAG Agents, and Structure-Augmented RAG Agents.
(1) Simple RAG Agents
We implement a BM25 (Robertson & Walker, 1994) retriever as a strong lexical baseline: it scores documents by term frequency with saturation and inverse document frequency, with length normalization controlled by parameters and . BM25 remains competitive for exact-match queries and complements dense retrievers with robust precision on keyworded questions.
(2) Embedding-based RAG Agents
Contriever (Izacard et al., 2021) is an unsupervised dense retriever trained via contrastive learning on large text corpora, enabling semantic matching without labeled pairs. Text-Embedding-3-Small/Large (OpenAI, 2024) are OpenAI’s general-purpose embedding models offering a cost–quality trade-off (e.g., 1,536 vs. 3,072 dimensions) for search and retrieval. Qwen3-Embedding-4B (Zhang et al., 2025) is a 4B-parameter embedding/reranking model family geared toward multilingual retrieval and long-text understanding.
(3) Structure-Augmented RAG Agents
RAPTOR (Sarthi et al., 2024) is method building a hierarchical tree of recursive summaries (bottom-up clustering and abstraction) and retrieves across levels for long-document QA. GraphRAG (Edge et al., 2024) extracts a knowledge graph and community hierarchy, then performs graph-aware retrieval and summarization. MemoRAG (Qian et al., 2025) introduces a dual-system pipeline with a light “global-memory” model to guide retrieval and a stronger model for final answers. HippoRAG-v2 (Gutiérrez et al., 2025) extends hippocampal-inspired retrieval to improve factual, sense-making, and associative memory tasks over standard RAG. We also evaluate three open-sourced memory agents: Mem0, Cognee and Zep. Mem0 (Chhikara et al., 2025) provides a persistent agent memory layer for storing/retrieving user-specific knowledge to enhance personalization. Cognee (Markovic et al., 2025) is an open-source memory engine that builds structured (graph-native) memories via ECL pipelines to power graph-aware RAG. Zep (Rasmussen et al., 2025) is a temporal knowledge-graph memory platform for agents, designed to assemble and retrieve long-term conversational and business context.
B.3 Agentic Memory Agents
For Agentic Memory Agents, We evaluate the Self-RAG (Asai et al., 2023), MemGPT (Packer et al., 2023), and MIRIX (Wang & Chen, 2025) on our benchmark. Self-RAG use LLMs as the agent to decide when/what to retrieve and to critique its own outputs. MemGPT operates the hierarchical memory management, paging relevant snippets between short-term and long-term stores and using event-driven interrupts to maintain coherence and evolvability over extended interactions. MIRIX adopts a multi-agent memory architecture with six specialized memory types (Core, Episodic, Semantic, Procedural, Resource, Knowledge Vault) and a coordinator that orchestrates updates/retrieval across agents.
For comparability, we standardize prompts, tool access, and settings (like retrieval TopK and input chunk size) across above all systems.
Appendix C Prompts
We introduce the examples of prompt used memory construction and task execution in this section.
C.1 Instructions for Memory Construction
@@
C.2 Instructions for Task Execution
In Figure 5, we provide the examples of instructions used on different of datasets when handling the input queries. For some existing datasets, we refer the prompt settings from previous work such as (Hsieh et al., 2024; Wu et al., 2025). For the dataset Bench-Sum, we also insert two answer examples as demo in the prompt to help the agent better understand the questions and standardize its outputs.
Appendix D Detailed Experimental Results
@@
[…]
| GPT-4o | 96.0 | 96.0 | 90.0 | 87.0 | 69.0 |
| GPT-4o-mini | 93.0 | 93.0 | 87.0 | 73.0 | 66.0 |
| GPT-4.1-mini | 93.0 | 82.0 | 85.0 | 68.0 | 50.0 |
| Gemini-2.0-Flash | 91.0 | 90.0 | 84.0 | 88.0 | 67.0 |
| Claude-3.7-Sonnet | 97.0 | 98.0 | 86.0 | 87.0 | 79.0 |
[…]
| BM25 | 89.0 | 89.0 | 84.0 | 62.0 | 53.0 |
[…]
| Contriever | 89.0 | 88.0 | 80.0 | 55.0 | 41.0 |
| Text-Embed-3-Small | 88.0 | 89.0 | 83.0 | 54.0 | 36.0 |
| Text-Embed-3-Large | 90.0 | 91.0 | 80.0 | 55.0 | 46.0 |
| Qwen3-Embedding-4B | 90.0 | 88.0 | 86.0 | 67.0 | 59.0 |
[…]
| RAPTOR | 78.0 | 75.0 | 73.0 | 48.0 | 23.0 |
| GraphRAG | 64.0 | 54.0 | 49.0 | 24.0 | 6.0 |
| MemoRAG | 90.0 | 87.0 | 86.0 | 66.0 | 56.0 |
| HippoRAG-v2 | 81.0 | 86.0 | 73.0 | 38.0 | 29.0 |
| Mem0 | 35.0 | 37.0 | 35.0 | 29.0 | 26.0 |
| Cognee | 34.0 | 42.0 | 42.0 | 41.0 | 18.0 |
| Zep | 83.0 | 74.0 | 70.0 | 50.0 | 37.0 |
[…]
| Self-RAG | 19.0 | 13.0 | 6.0 | 15.0 | 5.0 |
| MemGPT | 89.0 | 83.0 | 79.0 | 56.0 | 31.0 |
| MIRIX | 42.0 | 53.0 | 49.0 | 36.0 | 12.0 |
| MIRIX(4.1-mini) | 65.0 | 83.0 | 69.0 | 73.0 | 25.0 |
D.1 Detailed Results on TTL
We give detailed results on Multi-Class Classification (MCC) task in Table D. For all three types of tasks, RAG-based agents generally underperform compared to their respective GPT-4o-mini backbones. This observation highlights certain limitations inherent to the RAG approach. For instance, in TTL tasks, RAG-based methods often struggle to more accurately retrieve context from memory that is closely associated with the input.
D.2 Results on Input Chunk Size Ablation Study
In Table 9, we report the detailed results on evaluating the RAG-based Agents on different chunk sizes and datasets. We selected chunk sizes from the two sets {512, 4096} and {512, 1024, 2048, 4096}.
D.3 Results on Retrieval TopK Ablation Study
In Table 9, we report the detailed results of the selected RAG-based Agents evaluated on five datasets. We choose different TopK ranging from {2, 5, 10}.
D.4 Results on Different Context Length Ablation Study